# Nested and outlier embeddings into trees

### v1

@@ 3.1 Bad metrics
To motivate this work, we may begin by asking: do there exist metrics for which removing a small number of points results in a significantly better probabilistic embedding into HSTs? In this subsection, we will discuss types of metrics for which this is the case. In particular, in Figure 1, we give an example of an unweighted graph metric that is formed by “linking” two unweighted graphs through an edge. If we select one of these graphs to be $K_{n-\log n}$ (the complete graph on $n-\log n$ nodes) and the other to be a constant degree expander graph on $\log n$ nodes (denoted $G_{\log n}$), we get a graph on $n$ total nodes. However, Bartal [11] shows that, for any $x$, certain expander graphs $G_{x}$ incur a distortion $\Omega(\log x)$ when probabilistically embedding into HSTs. Thus, the graph in Figure 1 requires $\Omega(\log\log n)$ distortion for HST embeddings (inheriting the distortion bound of the submetric $G_{\log n}$), whereas the same graph with $G_{\log n}$ removed can be embedded with $O(1)$ distortion.
$K_{n-\log n}$${\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}G_{\log n}}$
Figure 1: We can consider an unweighted graph on $n$ nodes such that one part of the graph is $K_{n-\log n}$ (the complete graph on $n-\log n$ nodes), and the remaining $\log n$ nodes form some constant-degree expander graph $G_{\log n}$. The two parts of the graph can then be connected by a single edge. Note that the subgraph induced by each of these two sub-parts of this graph is an isometric subgraph of the original graph.
@@
Definition 3.1 (Metric composition [13]):
@@
The highest distortion required to embed some metric in $\set{M}\cup\mathcal{N}$ into HSTs is a lower bound on the minimum distortion to embed $M_{\beta}[\mathcal{N}]$ into HSTs, and in Theorem 3.2 we show that in fact this quantity is an upper bound as well (up to constant factors).
Theorem 3.2:
Let $(M,d_{M})$ be a metric with distortion at most $c$ for probabilistically embedding into HSTs, and let the metrics in $\mathcal{N}=\set{(N_{x},d_{N_{x}})}_{x\in M}$ also probabilistically embed with distortion at most $c$. Then there is a probabilistic embedding of $M_{\beta}[\mathcal{N}]$ into HSTs with distortion at most $O(c)$.
The proof of the theorem is deferred to Appendix A.1.
3.2 Buy-at-Bulk
In the (uniform) buy-at-bulk problem, [^4] We assume the metric version of the problem rather than the graph version, as Awerbuch and Azar [6] note that they are equivalent. we are given a metric space $\mathcal{X}$ as well as a set of demand pairs $(s_{i},t_{i})$, each with a demand $d_{i}$ and a sub-additive cost function $g$. (We will assume wlog that $g(1)=1$, as this is always possible by rescaling.) The goal is to select a single path $P_{i}$ for each demand pair in a way that minimizes $\sum_{e}\ell_{e}g(\sum_{i:e\in P_{i}}d_{i})$, where $\ell_{e}$ is the length of edge $e$ and all the sum is over all edges. This has natural applications in many areas of operation research.
A key result of Awerbuch and Azar [6] states that if $\mathcal{D}$ is a probabilistic tree embedding of metric $(X,d)$ with distortion $c$, then the optimal cost of the buy-at-bulk problem on $(X,d)$ is at most a factor of $c$ less than the expected value of the optimal solution to the same problem on the embedded metric. The buy-at-bulk problem is trivial on trees, where there is only one path between each pair of nodes. Thus, using the results of Munagala et al. [39] we can obtain an expected approximation ratio $O(c)$, where $c$ is the optimal distortion for probabilistically embedding the input into HSTs.
@@ 3.2 Buy-at-Bulk
For a given metric $\mathcal{X}$ with weights $w$ and any choice of $c\geq 1$, we will define $\text{OutlierCost}(c):=\min_{K\subseteq X;(X\setminus K)\xhookrightarrow{c}_{p}\mathcal{H}}w(K)$ to be the cost of the cheapest set we can leave out while still embedding the remaining metric into HSTs with at most $c$ distortion. We will also be interested in the size of such a cheap set, so we define $\mathcal{K}(c):=|\arg\min_{K\subseteq X;(X\setminus K)\xhookrightarrow{c}_{p}\mathcal{H}}w(K)|$. In Theorem 3.3, we give a bound on the cost of Algorithm 1. Note that for some constant $\eta$, if $c=\eta\log n$, [^6] In this section, we let $\eta$ be a constant such that for all $n$-point metric spaces $X$, $X\xhookrightarrow{\eta\log n}_{p}\mathcal{H}$. (Such a constant is guaranteed to exist by [28].) we have $\text{OutlierCost}(c)=0$, so Algorithm 1 never obtains worse than an $O(\log n)$ approximation.
@@
$O({c\cdot OPT+\text{OutlierCost}(c)\cdot\log^{2}(\mathcal{K}(c))}$.
We defer the formal proof to Appendix A.2.
@@
Now consider the result of running Algorithm 1 on this instance. Note that if we remove from $M^{*}$ the nodes corresponding to $N_{x^{*}}^{\prime}$, we get a metric that embeds into HSTs with constant distortion, by Theorem 3.2. We also know that the outlier cost of the nodes corresponding to $N_{x^{*}}^{\prime}$ is at most $\log^{2}n$, as there are $O(\log n)$ demand pairs associated with these nodes; each such pair has demand $1$; nodes in $N_{x^{*}}^{\prime}$ only participate in demand pairs with other nodes in $N_{x^{*}}^{\prime}$; and the diameter of $N_{x^{*}}^{\prime}$ is at most $\log n$ (due to being the path metric of an unweighted graph on $\log n$ nodes). Thus, Algorithm 1 gives us a solution of cost at most $O(OPT+\log^{4}n)=O(OPT)$, where we used the fact that $OPT\geq\ell\geq\log^{4}n$, as there is some demand pair with distance at least $\ell$.
3.3 Dial-A-Ride
@@
However, we will need to use a more complicated algorithm $\mathcal{A}$ to solve the dial-a-ride problem on HSTs than we did for buy-at-bulk. In particular, we will consider an algorithm $\mathcal{A}$ that gets approximation ratio $r_{n,m,\ell}$ for dial-a-ride instances on HSTs with $n$ nodes, $m$ requests, and $\ell$ van capacity. In Theorem 3.4, we consider the approximation ratio achieved by Algorithm 1 with such an input $\mathcal{A}$. Again note that for $c=\eta\cdot\log n$, $\text{OutlierCost}(c)=0$, so the worst-case approximation ratio is $O(\log n)$.
Theorem 3.4:
@@
Let $(\mathcal{X},\set{(s_{i},t_{i})}_{i=1}^{m},\ell)$ be a dial-a-ride input such that $OPT$ is the cost of an optimal solution for this instance, $\Delta$ is the diameter of $\mathcal{X}$, and $\text{OutlierCost}(c)$ and $\mathcal{K}(c)$ are as in Section 3.2. Then Algorithm 1 achieves a feasible solution with expected cost at most
@@
We prove this theorem in Appendix A.3. Note that this algorithm’s performance is generally better on instances whose optimal cost is much bigger than $\Delta$ (such as instances in which the total number of requests is large compared to $\ell\Delta$).

### v2

@@ 0.E.2 Some metrics have small outlier sets
To motivate this work, we may ask: do there exist metrics for which removing a small number of points results in a significantly better probabilistic embedding into HSTs? In this subsection, we discuss types of metrics for which this is the case. In particular, consider the distance metric on an unweighted graph obtained by “linking” two unweighted graphs through an edge, the first graph being an expander on $n$ nodes and the second being the complete graph on $n-\log n$ nodes. Bartal [10] shows that, for any $x$, certain expander graphs on $x$ nodes incur a distortion $\Omega(\log x)$ when probabilistically embedding into HSTs. Thus, the graph we have described requires $\Omega(\log\log n)$ distortion for HST embeddings (inheriting the distortion bound of the expander subgraph), whereas the same graph with the expander removed can be embedded with $O(1)$ distortion.
@@
Definition 8(Metric composition [12])
@@
The highest distortion required to embed a metric in $\set{M}\cup\mathcal{N}$ into HSTs is a lower bound on the minimum distortion to embed $M_{\beta}[\mathcal{N}]$ into HSTs, and careful composition reveals that in fact this quantity is an upper bound as well (up to constant factors).
0.E.3 Buy-at-bulk
In the (uniform) buy-at-bulk problem, [^6] We assume the metric version of the problem rather than the graph version, as Awerbuch et al. [6] note that they are equivalent. we are given a metric space $\mathcal{X}$ as well as a set of demand pairs $(s_{i},t_{i})$, each with a demand $d_{i}$ and a sub-additive cost function $g$. (We will assume wlog that $g(1)=1$, as this is always possible by rescaling.) The goal is to select a single path $P_{i}$ for each demand pair in a way that minimizes $\sum_{e}\ell_{e}g(\sum_{i:e\in P_{i}}d_{i})$, where $\ell_{e}$ is the length of edge $e$ and all the sum is over all edges. This has natural applications in many areas of operation research.
A key result of Awerbuch et al. [6] states that if $\mathcal{D}$ is a probabilistic tree embedding of metric $(X,d)$ with distortion $c$, then the optimal cost of the buy-at-bulk problem on $(X,d)$ is at most a factor of $c$ less than the expected value of the optimal solution to the same problem on the embedded metric. The buy-at-bulk problem is trivial on trees, where there is only one path between each pair of nodes. Thus, using the results of Munagala et al. [31] we can obtain an expected approximation ratio $O(c)$, where $c$ is the optimal distortion for probabilistically embedding the input into HSTs.
@@ 0.E.3 Buy-at-bulk
For a given metric $\mathcal{X}$ with weights $w$ and any choice of $c\geq 1$, we will define $\text{OutlierCost}(c):=\min_{K\subseteq X;(X\setminus K)\xhookrightarrow{c}_{p}\mathcal{H}}w(K)$ to be the cost of the cheapest set we can leave out while still embedding the remaining metric into HSTs with at most $c$ distortion. We will also be interested in the size of such a cheap set, so we define $\mathcal{K}(c):=|\arg\min_{K\subseteq X;(X\setminus K)\xhookrightarrow{c}_{p}\mathcal{H}}w(K)|$. In Claim 3, we give a bound on the cost of Algorithm 6. Note that for some constant $\eta$, if $c=\eta\log n$, [^8] In this section, we let $\eta$ be a constant such that for all $n$-point metric spaces $X$, $X\xhookrightarrow{\eta\log n}_{p}\mathcal{H}$. (Such a constant is guaranteed to exist by [23].) we have $\text{OutlierCost}(c)=0$, so Algorithm 6 never obtains worse than an $O(\log n)$ approximation.
@@
In fact, for any $c$ directly considered by the algorithm in line 6, the algorithm’s cost is bounded by $O({c\cdot OPT+\text{OutlierCost}(c)\cdot\log^{2}(\mathcal{K}(c))}$.
First we briefly justify why proving the second part of the statement is sufficient to prove the overall statement, which minimizes over all $c$ not just the $c_{i}$ considered by the algorithm. First, for all $c\geq\eta\log n=c_{0}$, $\text{OutlierCost}(c)=0$, so we need only justify this for $c<\eta\log n$. Let $c$ be the value that minimizes $c\cdot OPT+\text{OutlierCost}(c)\log^{2}(n)$, and note that any $c_{i}>c$ has optimal outlier set value at most equal to that of $c$. As the algorithm will consider some value of $c_{i}$ that is bigger than $c$ and at most $\eta$ times its value (if $c\geq\eta/2$, it will in fact consider a choice of $c_{i}$ that is at most twice this value), we will end up with the desired approximation.
Proof of Claim 3. We show that if index $i$ is selected as the optimal solution by the algorithm, then the cost of the solution is at most $O(c_{i}\cdot OPT+\text{OutlierCost}(c_{i})\log^{2}(\mathcal{K}(c_{i})))$. Let $K_{i}$ be the outlier set from line 8 of Algorithm 6.
@@
Now consider the result of running Algorithm 6 on this instance. Note that if we remove from $M^{*}$ the nodes corresponding to $N_{x^{*}}^{\prime}$, we get a metric that embeds into HSTs with constant distortion. We also know that the outlier cost of the nodes corresponding to $N_{x^{*}}^{\prime}$ is at most $\log^{2}n$, as there are $O(\log n)$ demand pairs associated with these nodes; each such pair has demand $1$; nodes in $N_{x^{*}}^{\prime}$ only participate in demand pairs with other nodes in $N_{x^{*}}^{\prime}$; and the diameter of $N_{x^{*}}^{\prime}$ is at most $\log n$ (due to being the path metric of an unweighted graph on $\log n$ nodes). Thus, Algorithm 6 gives us a solution of cost at most $O(OPT+\log^{4}n)=O(OPT)$, where we used the fact that $OPT\geq\ell\geq\log^{4}n$, as there is some demand pair with distance at least $\ell$.
0.E.4 Dial-a-ride
@@
However, we will need to use a more complicated algorithm $\mathcal{A}$ to solve the dial-a-ride problem on HSTs than we did for buy-at-bulk. In particular, we will consider an algorithm $\mathcal{A}$ that gets approximation ratio $r_{n,m,\ell}$ for dial-a-ride instances on HSTs with $n$ nodes, $m$ requests, and $\ell$ van capacity. In Claim 4, we consider the approximation ratio achieved by Algorithm 6 with such an input $\mathcal{A}$. Again note that for $c=\eta\cdot\log n$, $\text{OutlierCost}(c)=0$, so the worst-case approximation ratio is $O(\log n)$.
Claim 4
@@
Let $(\mathcal{X},\set{(s_{i},t_{i})}_{i=1}^{m},\ell)$ be a dial-a-ride input such that $OPT$ is the cost of an optimal solution for this instance, $\Delta$ is the diameter of $\mathcal{X}$, and $\text{OutlierCost}(c)$ and $\mathcal{K}(c)$ are as in Section 0.E.3. Then Algorithm 6 achieves a feasible solution with expected cost at most
@@
Note that this algorithm’s performance is generally better on instances whose optimal cost is much bigger than $\Delta$ (such as instances in which the total number of requests is large compared to $\ell\Delta$).